OpenStudy (campayne):

Use substitution to determine if the given ordered pair is an element of the solution set for the system of equations. Type solution or not a solution. (5, 3);{6x - 7y = 1 3x + 7y = 5

563blackghost (563blackghost):

Well a coordinate is in the format \(\large\bf{(x,y)}\) so we plug `5 as x` and `3 as y`.... \(\huge\bf{6(5)-7(3)=1}\) \(\huge\bf{3(5) + 7(3)=5}\) Are they each `true` to their total?

OpenStudy (campayne):

seems like it

Directrix (directrix):

Calculate this: (5, 3) 6x - 7y = 1 What is 6 times 5 = What is -7 * 3 = Add those two and see if you get an answer of 1.

OpenStudy (campayne):

and if i do

Directrix (directrix):

If you do, that means that (5, 3) is on graph of 6x - 7y = 1. The next step is to follow a similar procedure to determine is (5,3) is also a solution of 3x + 7y = 5.